Read the problem below and find the solution. Use a model or act the problem out to help solve it.
The Boxerville Fighting Hens's record has improved every year. The 2000 team won 3 more games than the 1999 team, and the 2001 team won 2 more games than the 2000 team. If the 2001 team won 20 games, how many did the 1999 team win?


Answer: ????

Answer by Theo(13342)

1999 = x
2000 = x + 3
2001 = x + 3 + 2 = x + 5
2001 = 20
x + 5 = 20
x = 15
1999 = 15
2000 = 15 = 3 = 18
2001 = 15 + 5 = 20
your answer is that the 1999 team won 15 games.